# SENTRA: Selected-Next-Token Transformer for LLM Text Detection

### v1

@@ 4.4 Results
[…]
@@
Table 3 shows that SENTRA outperformed all the baselines on average and in the worst case across the three OOD and one OOLLM evaluations. SENTRA achieved average AUROC performance improvements of 1.8%, 5.4% and 6.7% for RAID [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch], M4GT [Wang et al.(2024a)Wang, Mansurov, Ivanov, Su, Shelmanov, Tsvigun, Mohammed Afzal, Mahmoud, Puccetti, Arnold, Aji, Habash, Gurevych, and Nakov] and MAGE [Li et al.(2024a)Li, Li, Cui, Bi, Wang, Wang, Yang, Shi, and Zhang] out-of-domain datasets respectively, as compared to the second-best performing baseline. For the OOLLM evaluation, SENTRA showed a 7.5% increase over the next best baseline. These results show SENTRA serves as a generalizable encoder for LLM detection models when one considers likely OOD or OOLLM distribution shifts. These results show, in the likely event your detector encounters a domain outside the training distribution, we expect SENTRA to have the best expected performance and best worst-case performance on those unseen domains.
[…]
@@ 4.5 LLM Prompting Case Study
To control inference costs, we limited the evaluation to 100 samples per domain/model, using the same datasets from OOD and OOLLM experiments. The evaluation results are presented in Table 4, alongside SENTRA’s performance. Overall, the reasoning-based o1 model demonstrated stronger detection capabilities than the standard 4o model, particularly on RAID-OOD and M4GT-OOD. Nevertheless, SENTRA consistently outperformed both OpenAI models across all datasets.
@@ 5 Conclusions
In this paper, we proposed a novel general purpose supervised LLM text detector method SENTRA that is a Transformer-based encoder leveraging SNTP sequences and utilizing contrastive pre-training on large amounts of unlabeled data. We show this supervised method acting on SNTP input outperforms previously considered heuristic functions and other methods that rely on text input. Since supervised detectors tend to perform better on data that is similar to their training distributions [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch], it is essential to include a wide variety of domains when testing such general-purpose detectors. Therefore, we tested the performance of SENTRA on three public datasets RAID, M4GT and MAGE containing a broad range of different domains (24 in total) across various experimental settings and compared its performance with eight popular baselines. We also evaluated SENTRA and the baselines on a out-of-LLM evaluation.
@@ 6 Limitations
In this work, we studied the effects of domain shifts on detection models. While these have significant impacts on detector performance, other factors can also influence results. Notably, prompt variation can have a large effect on detectors [Kumarage et al.(2023b)Kumarage, Sheth, Moraffah, Garland, and Liu]. Many LLM detection benchmark datasets use prompt templates [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch] to generate their samples. However, these templates exhibit significantly less prompt variety than what a real-world detector is likely to encounter. Benchmark datasets with a broader range of prompting strategies are needed to further assess the robustness of detection methods.
We pre-trained our model on a relatively small sample of Common Crawl data. The volume of data and the amount of compute used for pre-training were small relative to what is typically used for foundation models [Liu et al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis, Zettlemoyer, and Stoyanov, Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever]. It is very likely SENTRA could be significantly improved with additional pre-training on larger datasets.
@@ Appendix B Additional Results and Experimental Notes
Here we show the mean and standard deviation across three runs, (random seeds 42,43,44) for the methods that are not zero shot or logistic regression based. Note there were three M4GT and four RAID samples where Ghostbuster could not make an inference due to the low number of tokens in the document. For this documents, we infilled a low prediction score indicating human prediction. For the RAID dataset, we used the Binoculars for each document released by [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch].
@@ Appendix C Computational Complexity
LLM generators are computationally expensive. Unfortunately, methods that rely on SNTP inputs depend on LLM inference, making it the most costly component of all detection methods studied in this work. However, SENTRA is a relatively small model with only eight Transformer layers, meaning that computational costs at inference are dominated by the production of SNTP inputs. During training, we cache the SNTP sequences so that the LLMs are run only once per sample. SENTRA uses the same LLMs as Binoculars [Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein], and because the cost of the SENTRA encoder is minimal compared to LLM inference, the overall computational complexity of SENTRA is roughly equivalent to that of the Binoculars method. Refer to Table 14 for detailed number of parameters.

### v2

@@ 4.4 Results
| RoBERTa [21] | 90.9 | 84.4 | 88.2 | 82.8 | 88.3 | 74.4 | 87.1 | 69.9 |
| Text-Fluoroscopy [43] | 76.4 | 70.6 | 83.2 | 78.1 | 63.9 | 47.8 | 41.5 | 28.3 |
| UAR-D [33] | 81.7 | 71.4 | 75.3 | 63.9 | 63.4 | 40.5 | 71.7 | 65.8 |
| UAR-L [33] | 87.3 | 76.3 | 84.7 | 71.0 | 76.4 | 61.2 | 80.4 | 70.7 |
| Ghostbuster [38] | 84.7 | 74.1 | 87.8 | 73.3 | 79.2 | 65.0 | 68.5 | 34.3 |
| PPL [9] | 72.9 | 69.4 | 87.0 | 81.7 | 57.2 | 45.7 | 59.0 | 25.4 |
| Binoculars [11] | 82.0 | 79.4 | 89.1 | 79.0 | 61.7 | 52.9 | 61.8 | 14.7 |
| Fast-DetectGPT [2] | 78.6 | 75.6 | 87.4 | 79.1 | 63.0 | 54.9 | 37.9 | 2.8 |
@@
Table 3 shows that SENTRA outperformed all the baselines on average and in the worst case across the three OOD and one OOLLM evaluations. SENTRA achieved average AUROC performance improvements of 1.8%, 5.4% and 6.7% for RAID Dugan et al. (2024), M4GT Wang et al. (2024a) and MAGE Li et al. (2024a) out-of-domain datasets respectively, as compared to the second-best performing baseline. For the OOLLM evaluation, SENTRA showed a 7.5% increase over the next best baseline. These results show SENTRA serves as a generalizable encoder for LLM detection models when one considers likely OOD or OOLLM distribution shifts. These results show, in the likely event your detector encounters a domain outside the training distribution, we expect SENTRA to have the best expected performance and best worst-case performance on those unseen domains.
Since LLMs became increasingly available and their usage has surged, interest in detection tools, such as those presented in this paper, has grown Wu et al. (2023). At the same time, countermeasures have emerged to attack these LLM text detectors, typically by altering LLM-generated text to elicit false negatives Koike et al. (2024). Dugan et al. (2024) demonstrated many attacks can significantly degrade detector performance. In that study, the best open-source tool, Binoculars Hans et al. (2024), exhibited much stronger performance on non-attacked data than on attacked data. For the unsupervised methods, Guo et al. (2023); Hans et al. (2024); Bao et al. (2024), it is not immediately clear how to adapt the approach to a known attack. In contrast, for the supervised methods, the adaptation strategy is straightforward: train on attacked data. A model that is robust to a known attack, like the common paraphrase attack, should be able to detect LLM generated text even if that attack appears in a new domain. The RAID-OOD Dugan et al. (2024) dataset demonstrates this situation where 11 attacks appear in the training and test sets. The results in Table 3 show SENTRA outperformed other methods when training and evaluating in the out-of-domain scenario where known attacks are included.
@@ 4.5 LLM Prompting Case Study
To control inference costs, we limited the evaluation to 100 samples per domain/model, using the same datasets from the OOD and OOLLM experiments. The evaluation results are presented in Table 4, alongside SENTRA’s performance. Overall, the reasoning-based o1 model demonstrated stronger detection capabilities than the standard 4o model, particularly on RAID-OOD and M4GT-OOD. Nevertheless, SENTRA consistently outperformed both OpenAI models across all datasets.
@@ 5 Conclusions
In this paper, we proposed a novel general purpose supervised LLM text detector method SENTRA that is a Transformer-based encoder leveraging SNTP sequences and utilizing contrastive pre-training on large amounts of unlabeled data. We show this supervised method acting on SNTP input outperforms previously considered heuristic functions and other methods that rely on text input. Since supervised detectors tend to perform better on data that is similar to their training distributions Dugan et al. (2024), it is essential to include a wide variety of domains when testing such general-purpose detectors. Therefore, we tested the performance of SENTRA on three public datasets RAID, M4GT and MAGE containing a broad range of different domains (24 in total) across various experimental settings and compared its performance with eight popular baselines. We also evaluated SENTRA and the baselines on a out-of-LLM evaluation.
@@ 6 Limitations
In this work, we studied the effects of domain shifts on detection models. While these have significant impacts on detector performance, other factors can also influence results. Notably, prompt variation can have a large effect on detectors Kumarage et al. (2023b). Many LLM detection benchmark datasets use prompt templates Dugan et al. (2024) to generate their samples. However, these templates exhibit significantly less prompt variety than what a real-world detector is likely to encounter. Benchmark datasets with a broader range of prompting strategies are needed to further assess the robustness of detection methods.
We pre-trained our model on a relatively small sample of Common Crawl data. The volume of data and the amount of compute used for pre-training were small relative to what is typically used for foundation models Liu et al. (2019); Radford et al. (2021). It is very likely SENTRA could be significantly improved with additional pre-training on larger datasets.
@@ Appendix B Additional Results and Experimental Notes
Here we show the mean and standard deviation across three runs, (random seeds 42,43,44) for the methods that are not zero shot or logistic regression based. Note there were three M4GT and four RAID samples where Ghostbuster could not make an inference due to the low number of tokens in the document. For this documents, we infilled a low prediction score indicating human prediction. For the RAID dataset, we used the Binoculars for each document released by Dugan et al. (2024).
@@ Appendix C Computational Complexity
LLM generators are computationally expensive. Unfortunately, methods that rely on SNTP inputs depend on LLM inference, making it the most costly component of all detection methods studied in this work. However, SENTRA is a relatively small model with only eight Transformer layers, meaning that computational costs at inference are dominated by the production of SNTP inputs. During training, we cache the SNTP sequences so that the LLMs are run only once per sample. SENTRA uses the same LLMs as Binoculars Hans et al. (2024), and because the cost of the SENTRA encoder is minimal compared to LLM inference, the overall computational complexity of SENTRA is roughly equivalent to that of the Binoculars method. Refer to Table 14 for detailed number of parameters.